Suppose that , firm under perfectly competition market produce two commodities X1 and X2 with corresponding prices birr 10 and birr 15 . If cost function of the firm is TC= 2x12 +x1x2+2x22 where x1 and x2 denote the level of output, then, determine the following questions.

i. Profit maximizing level of output x1 and x2

ii. The amount of maximum profit


1
Expert's answer
2021-12-20T10:21:17-0500

Price of"\space X_1 =P_1=10"

Price of "X _\n2\n\u200b\n =P_ \n2\n\u200b\n =15"

Cost function "C=2X_1^2+X_1X_2+2X_2^2"

"Total\space Revenue\space R=P_1X_1+P_2X_2\\=10X_1+15X_2"

"Profit\space \pi=R-C\\=10X_1+15X_2-2X_1^2-X_1X_2-2X_2^2"

To find the profit maximizing output of "X_1:"

"\frac{\delta \pi}{\delta X_1}=0"

"\implies\frac{\delta \pi}{\delta X_1}(10X_1+15X_2-2X_1^2-X_1X_2-2X_2^2)=0"

"\implies10+0-4X_1-X_2-0=0"

"\implies4X_1+X_2=10.........................................1"


To find the profit maximizing output of "X_2:"

"\frac{\delta \pi}{\delta X_2}=0"

"\implies\frac{\delta \pi}{\delta X_2}(10X_1+15X_2-2X_1^2-X_1X_2-2X_2^2)=0"

"\implies0+15-0-X_1-4X_2=0"

"\implies X_1+4X_2=15...........................................2"


"Eq1\times4\implies16X_1+4X_2=40\\\underline{Eq2\times1\implies X_1\space \space \space \space +4X_2=15}\\\space \space \space \space \space \space\space\space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space \space 15X_1\space \space \space \space \space \space\space \space \space \space \space \space =25\\X_1=\frac{25}{15}\\\implies X=\frac{g5}{3}\\X_1\approx1\space unit"


In Eq

1, "4X_1+X_2=10\\\implies X_2=10-4(\frac{5}{3})\\X_2=\frac{30-20}{3}\\X_2=(\frac{10}{3})\\X_2\approx3 \space units"